In sampling for cresol (MW 108), 4.5 liters of workplace air from a copper smelting plant was collected in a Mylar bag. The worksite air temperature was 95oF and the atmospheric pressure was 740 mmHg. Upon analysis at 71oF laboratory with an atmospheric pressure of 765 mmHg, the concentration was found to be 348 mg/m3. What was the concentration at the worksite in ppm?
Answer #1

Lab conditions P = 765 mmHg = ( 765 / 760) atm = 1.007 atm

T = 71 F = ( 71-32) / 1.8 C = 217. C = 21.7 + 273K = 295 K

Under starting conditions P = 740 mmHg = ( 740/760) = 0.9737 atm ,  

T = 95 F = ( 95-32) / 1.8 = 35 C = 35 + 273 K = 308 K

we had formula PV = nRT

( n/V) = P/RT ( where n/V indicates mol / volume which we can say as concentration )

now ( concentration in lab / ( concentration at starting = ( P/T ) lab / ( P/T starting) = ( Plab x T starting) / ( P start x T lab)

348 mg/m3 / ( conc starting) = ( 1.007 atm x 308 K) / ( 0.9737 atm x 295K )

conc starting = 322 mg/m3

= ( 322 mg/m3) x ( 1m3 / 1000L) = 0.322 mg/L = ( 0.322 mg/L) x ( 1g/1000mg)

= 0.000322 g /L = 0.000322 g/L x ( 1L / 1000 ml) = 0.322 x 10^ -6 g/ml

W have  10^ -6 g/ml is equal to 1ppm

hence concentration in ppm = 0.322 ppm